# thanks: deceasedthanks: deceased

### v1

@@ Abstract
We conducted measurements of $K^{-}+{{}^{3}{\rm He}}\to\pi Y\!N+N^{\prime}$ reactions using $1~{}{\rm GeV}/c$ $K^{-}$-beam, aiming to understand the broad decay width of $\bar{K}\!N\!N$ ($\approx$ twice as broad as that of $\Lambda(1405)$ considered to be the $\bar{K}\!N$ quasi-bound state). We successfully reproduced distributions of the $\pi Y\!N$ invariant mass and momentum transfer of $\pi Y\!N$ by employing model fitting functions for $\bar{K}\!N\!N$ formation and quasi-free $\bar{K}$ absorption (${\rm QF}_{\bar{K}-{\rm abs}}$) processes. The model can describe the experimental data quite well, and four $\bar{K}NN\to\pi Y\!N$ cross-sections were obtained. The result suggests that the mesonic decay is the dominant decay branch of $\bar{K}\!N\!N$. It was also suggested that $\Gamma_{\pi\Lambda N}\sim\Gamma_{\pi\Sigma N}$, which indicates a significant contribution of the $I_{\bar{K}\!N}=1$ absorption channel to the $\bar{K}\!N\!N$ decay as well as the $I_{\bar{K}\!N}=0$ makes the $\bar{K}\!N\!N$ state $\approx$ twice as unstable as $\Lambda$(1405).
@@ I Introduction
Quasi-bound systems of antikaon ($\bar{K}$) and nucleus, so-called kaonic nuclei, represents an exotic nuclear state wherein a meson exists as a constituent particle inside nuclei. The existence of a kaonic nucleus is predicted as a natural extension of the interpretation of the $\Lambda(1405)$, which is not considered a simple three-quark state but rather a $\bar{K}$-$N$ hadronic molecule attributed to the strongly attractive $\bar{K}N$ interaction in the isospin $I_{\bar{K}N}=0$ channel. Properties of kaonic nuclei, such as binding energies, widths, and branching ratios, nor even confirmation of the existence of kaonic nuclei, are essential to construct a precise model of the $\bar{K}N$ interaction. Understanding these properties allows us to determine the energy dependence of the $\bar{K}\!N$ interaction or the $\bar{K}$-absorption strength to nucleon(s) in the energy region below the $\bar{K}N$ mass threshold. This energy region cannot be directly probed by $\bar{K}N$ scattering or $K^{-}$-atom experiments.
@@
Among kaonic nuclei, the lightest system, $\bar{K}\!N\!N$, considered to have quantum numbers $I=1/2$ and $J^{P}=0^{-}$, is essential for understanding the system’s fundamental features. We conducted the J-PARC E15 experiment to search for the $\bar{K}\!N\!N$ and reported the observation of the $I_{3}=+1/2~{}\bar{K}\!N\!N$ state, also known as ”$K^{-}pp$”, produced in the ${}^{3}{\rm He}(K^{-},~{}n)$ reactions and decaying into the non-mesonic $\Lambda p$ decay channel [1, 2]. The binding energy and decay width of $\bar{K}\!N\!N^{(I_{3}=+1/2)}$ were determined to be approximately $B_{\bar{K}\!N\!N}\sim 40~{}{\rm MeV}$ and $\Gamma_{\bar{K}\!N\!N}\sim 100~{}{\rm MeV}$, respectively, as the Breit-Wigner parameters.
@@
In theoretical studies, the binding energy and decay width of $\bar{K}\!N\!N$ are calculated to be in the range of $10-50~{}{\rm MeV}$ and $40-70~{}{\rm MeV}$, respectively, using the chiral-SU(3) based potential [3, 4, 5, 6, 7, 8, 9]. Theoretical studies with phenomenological potentials reported significantly larger binding energies by $\sim 20~{}{\rm MeV}$ and similar decay widths [5, 6, 7, 8, 9]. Therefore, the obtained $B_{\bar{K}\!N\!N}$ is within the range of the theoretical calculations. However, $\Gamma_{\bar{K}\!N\!N}$ is larger than those predictions. Since the $\Lambda(1405)$ resonance is considered as the $\bar{K}\!N$ quasi-bound state, the dominant decay branch of $\bar{K}\!N\!N$ is naively expected to be $\bar{K}\!N\!N\to\pi\Sigma N$, similar to the $\Lambda(1405)\to\pi\Sigma$ decay (100%). Then $\Gamma_{\bar{K}\!N\!N}$ would be expected to be similar to the decay width of $\Lambda(1405)$. However, the obtained $\Gamma_{\bar{K}\!N\!N}$ is approximately twice as large as $\Gamma_{\Lambda(1405)}$.
@@
As a reason for the large $\Gamma_{\bar{K}\!N\!N}$, two reasonable scenarios exist. One is that the $\bar{K}\!N\!N$ may decay into non-mesonic $Y\!N$ channels with similar strength to the $\pi\Sigma N$ channel. This non-mesonic decay occurs via $\bar{K}$ absorption by two nucleons in the nucleus, which most theoretical calculations did not consider. Another possibility is that the $I_{\bar{K}\!N}=1$ interaction may play an important role, similar to the $I_{\bar{K}\!N}=0$ interaction, at least in the absorption channels, i.e. $\bar{K}\!N\!N\to[\pi Y]_{I=1}N$, in contrast to the theoretical predictions [10, 11]. These two decay modes cannot occur in the decay of $\Lambda(1405)$ and newly open up in the $\bar{K}\!N\!N$ three-body system. Therefore, studying the mesonic decay modes is essential to clarify why $\Gamma_{\bar{K}\!N\!N}$ is much larger than $\Lambda(1405)$.
@@
To investigate the mesonic decay branches of $\bar{K}\!N\!N$, we analyzed $K^{-}+^{3}{\rm He}$ reactions resulting in $\pi Y\!N\!N$ mesonic final states using data from the J-PARC E15 experiment. The measured final states in this study are $\pi^{-}\Lambda pp$, $\pi^{+}\Lambda nn$, and $\pi^{\mp}\Sigma^{\pm}pn$. As described in Ref. [2], the $\bar{K}\!N\!N$ state is produced through the $\bar{K}$ and $N$ substitution via a nucleon knockout reaction on ${{}^{3}\rm He}$. Therefore, we focus on the $\pi Y\!N\!N$ channels that originate from the reaction processes:
@@
where $N^{\prime}$ is emitted in the forward direction. Subsequently, the backscattered intermediate $\bar{K}$ is absorbed by the two spectator nucleons as:
@@
where $\bar{K}\!N\!N$ production could contribute. The relevant $K^{-}+{{}^{3}\rm He}$ reaction channels of interest can be summarized as:
@@
where $\{\pi Y\!N\}$ is the particle set to be investigated, and $N^{\prime}$ is the emitted nucleon in the primary reaction (Eq.( 1)). Each final state is reconstructed as described on the right side of the equations, where $[\pi(\pi N)N]_{\rm det}$ represents detected particles, and $N_{\rm miss}$ is an undetected particle identified using the missing mass method. The $\bar{K}\!N\!N$ signal is expected to be observed in the $[\pi Y\!N]$ invariant-mass spectra for the $I_{3}=-1/2$ state in Eq.(3), and the $I_{3}=+1/2$ state in Eqs.(4) and (5). Following the analysis in Ref. [2], we performed a two-dimensional analysis on the $[\pi Y\!N]$ invariant-mass and the momentum transfer to the $[\pi Y\!N]$ system to isolate the signal of $\bar{K}\!N\!N$.
These reaction processes may couple to $Y^{*}$ resonances. In particular, two $Y^{*}$ resonances below the $\bar{K}\!N$ mass threshold, $\Lambda(1405)$ and $\Sigma(1385)$, would have an important contribution to the $\bar{K}$ absorption inside $\bar{K}\!N\!N$. Therefore, we analyzed the invariant-mass spectra of the $[\pi Y]$ subsystem to study the contributions of these $Y^{*}$ resonances.
This article is organized as follows. In Sec. II, we briefly introduce the J-PARC E15 experiment and describe the essential components of the detector system. In Sec. III, we present the analysis procedure used to select the $\pi Y\!N\!N$ channels. The obtained spectra will be shown in those results in Sec. IV. In Sec. V, we perform a model fitting using the functions introduced in Ref. [2]. Specifically, we discuss the $\bar{K}\!N\!N$ production and the quasi-free $\bar{K}$ absorption processes. In the fitting, we considered a background originating from the misidentification of $N^{\prime}$ and $N$ in Eqs.(1) and (2), respectively. Finally, we compare the cross sections of $K^{-}+{{}^{3}\rm He}$ reactions leading to $\pi Y\!N+N^{\prime}$ channels with the non-mesonic $\Lambda p+n$ channel.
@@ II Experiment
We briefly overview the experimental setup of J-PARC E15, focusing on the key apparatus used to analyze the $\pi Y\!N\!N$ channels. The experimental configuration has been described in detail in the references [12, 13, 14]. We took the current data set in parallel to the one used for the $K^{-}+\,^{3}{\rm He}\to\Lambda p+n$ analysis [2] with the same trigger condition. In the experiment, a focused $K^{-}$ beam interacted with a liquid 3He target. The particles produced in the $K^{-}+{{}^{3}\rm He}$ reaction were detected by a cylindrical detector system (CDS) surrounding the target. The CDS consists of a cylindrical drift chamber (CDC) and a cylindrical hodoscope counter made of a plastic scintillator (CDH).
The CDS operates in a magnetic field of $0.7~{}{\rm T}$ along the beam axis ($z$-axis). The tracking information provided by the CDC allows for the analysis of charged particle momenta. We determined the starting point of the charged particle or the reaction point from the CDC track and $K^{-}$ beam track by searching for the closest point of the two tracks. Additionally, particle identification for each track is performed using momentum and time-of-flight information obtained from the CDH.
We employed the CDH for the neutron detection, which consists of plastic scintillators measuring 30 mm in thickness and 790 mm in length. The inner surface of the CDH is positioned at a radius of 544 mm from the beam axis. Using a Monte Carlo simulation based on Geant4, we estimated the typical neutron detection efficiency to be 3–9% (at neutron momenta of 1–0.2 ${\rm GeV}/c$). The neutron momentum was determined using the time-of-flight information and neutron flight length obtained from the reaction point, and the $z$-position along the CDH deduced by the time difference between two PMTs at both ends.
@@ III Analysis
We focused on the following four channels: ${\pi^{-}\Lambda p}+p^{\prime}$ (Eq.(3)), ${\pi^{+}\Lambda n}+n^{\prime}$ (Eq.(4)), and ${\pi^{\mp}\Sigma^{\pm}p}+n^{\prime}$ (Eq.( 5)). Among these four channels, only the ${\pi^{-}\Lambda p}+p^{\prime}$ channel can be measured without detecting a neutron; namely, we required the detection of $(pp\pi^{-}\pi^{-})$ in the CDS. In the other three channels, neutron detection is necessary, along with detecting the remaining three charged particles ($p\pi^{+}\pi^{-}$).
The CDS detects the ${\pi Y\!N}$ particles in each final state. The undetected particle $N^{\prime}$ is identified through the missing mass method utilizing the measured $K^{-}$ beam and the information from the ${\pi Y\!N}$ particles. Other ${\pi Y\!N}+N^{\prime}$ channels, including multi-neutrons or $\gamma$ in the final decay products, were omitted from the present analysis since it is impossible to identify the final states without ambiguity.
@@ III.1 Neutron analysis
We applied the following criteria to identify neutrons. Firstly, we required no CDC track on the fired CDH segment and both sides to eliminate fake neutron signals from charged particles hitting the material close to the CDH segment. Secondly, we required the energy deposited by the CDH-hit to be greater than $2~{}{\rm MeV}ee$ (electron-equivalent) to remove accidental background recorded in the CDH. Finally, $1/\beta$ was calculated for the CDH-hit using the time-of-flight information, and events with $1.51/\beta<$ 1.5 were rejected to eliminate $\gamma$-ray background originating from the $K^{-}+{{{}^{3}}{\rm He}}$ reaction which peaks at $1/\beta\approx 1$.
Figure 1: Mass distributions of the missing particle $X$ in the ($pn\pi^{+}\pi^{-}+X$) event selection. (a) shows the distribution before background subtraction, and (b) shows the distribution after subtracting the background.
Even after implementing the background rejection procedures described above, our data still contains numerous neutron-like fake signals in the neutron’s missing mass spectrum. Figure 1-(a) illustrates the distribution of missing mass ($m_{X}$) in the $K^{-}+{{{}^{3}}{\rm He}}\to pn\pi^{+}\pi^{-}+X$ events. The figure shows that the signal-to-noise ratio is approximately one-to-one, necessitating a thorough evaluation of the background effect for detailed analysis.
The fake neutron signals are generated through multiple reactions or scattering, reach the CDH, and are identified as neutral particles. Therefore, we evaluated the impact of these fake signals by simulating random CDH hits using an event-mixing method. In this method, we disconnected CDH-hit information from other charged particle hit information and randomly recombined the two data sets. The gray hatched curve in the figure represents the result of applying this event-mixing method. We determined a scaling factor for the background distribution by fitting the $m_{X}$-distribution below $0.8~{}{\rm GeV}/c^{2}$, where the data should only include neutron contamination events.
Figure 1-(b) shows that the background subtraction is successful in the $m_{X}=m_{n}$ region, and the $[pn\pi^{+}\pi^{-}]+n_{\rm miss}$ final state remains clearly as a peak in the missing mass spectrum. This background subtraction procedure also agrees in the higher mass region up to 1.2 GeV/$c^{2}$, including the $m_{X}\sim m_{\Lambda}$ region. In fact, $\Lambda$ missing events, specifically $[pn\pi^{+}\pi^{-}]+\Lambda_{\rm miss}$ events, also remain clearly visible. This agreement demonstrates the validity of the fake-neutron subtraction procedure. We evaluated the fitting error of the scaling factor, as indicated by the gray box in the figure, and considered this error to be a systematic error in the subsequent analysis.

### v2

@@ Abstract
We conducted measurements of $K^{-}+{{}^{3}{\rm He}}\to\pi Y\!N+N^{\prime}$ reactions using a $1~{}{\rm GeV}/c$ $K^{-}$-beam, with the objective of understanding the broad decay width of $\bar{K}\!N\!N$ (approximately twice as broad as that of $\Lambda(1405)$ considered to be the $\bar{K}\!N$ quasi-bound state). We successfully reproduced distributions of the $\pi Y\!N$ invariant mass and momentum transfer for $\pi Y\!N$ using model fitting functions for $\bar{K}\!N\!N$ formation and quasi-free $\bar{K}$ absorption (${\rm QF}_{\bar{K}-{\rm abs}}$) processes. The model can describe the experimental data quite well, and four $\bar{K}\!N\!N\to\pi Y\!N$ cross-sections were obtained. The results indicate that mesonic decay is the dominant decay branch of $\bar{K}\!N\!N$. The results also suggest that $\Gamma_{\pi\Lambda N}\sim\Gamma_{\pi\Sigma N}$, which indicates that the $I_{\bar{K}\!N}=1$ absorption channel, in addition to the $I_{\bar{K}\!N}=0$ absorption channel, substantially contribute to the $\bar{K}\!N\!N$ decay, making the $\bar{K}\!N\!N$ state approximately twice as unstable as $\Lambda$(1405).
@@ I Introduction
Quasi-bound systems of an antikaon ($\bar{K}$) and nucleus—the so-called kaonic nuclei—represent an exotic nuclear state in which a meson exists as a constituent particle inside nuclei. The existence of a kaonic nucleus is predicted as a natural extension of the interpretation of the $\Lambda(1405)$, which is not considered a simple three-quark state but rather a $\bar{K}$-$N$ hadronic molecule attributable to the strongly attractive $\bar{K}\!N$ interaction in the isospin $I_{\bar{K}\!N}=0$ channel. Neither knowledge of the properties of kaonic nuclei, such as their binding energies, widths, and branching ratios, nor even confirmation of the existence of kaonic nuclei, is essential to construct a precise model of the $\bar{K}\!N$ interaction. Understanding these properties enables us to determine the energy dependence of the $\bar{K}\!N$ interaction or the $\bar{K}$-absorption strength to nucleon(s) in the energy region below the $\bar{K}\!N$ mass threshold. This energy region cannot be directly probed by $\bar{K}\!N$ scattering or $K^{-}$-atom experiments.
@@
Among kaonic nuclei, the lightest system, $\bar{K}\!N\!N$, which is considered to have quantum numbers $I=1/2$ and $J^{P}=0^{-}$, must be understood for the system’s fundamental features to be elucidated. We conducted the J-PARC E15 experiment to search for the $\bar{K}\!N\!N$ and reported observing the $I_{3}=+1/2~{}\bar{K}\!N\!N$ state, also known as ”$K^{-}pp$,” produced in the ${}^{3}{\rm He}(K^{-},~{}n)$ reactions and decaying into the non-mesonic $\Lambda p$ decay channel [1, 2]. The binding energy and decay width for $\bar{K}\!N\!N^{(I_{3}=+1/2)}$ were determined to be $B_{\bar{K}\!N\!N}\sim 40~{}{\rm MeV}$ and $\Gamma_{\bar{K}\!N\!N}\sim 100~{}{\rm MeV}$, respectively, as the Breit–Wigner parameters.
@@
In theoretical studies, the binding energy and decay width for $\bar{K}\!N\!N$ have been calculated to be in the range $10-50~{}{\rm MeV}$ and $40-70~{}{\rm MeV}$, respectively, using the chiral-SU(3)-based potential [3, 4, 5, 6, 7, 8, 9]. In theoretical studies with phenomenological potentials, substantially larger binding energies (by $\sim 20~{}{\rm MeV}$) and similar decay widths have been reported [5, 6, 7, 8, 9]. Therefore, the obtained $B_{\bar{K}\!N\!N}$ is within the range of the theoretical calculations. However, the obtained $\Gamma_{\bar{K}\!N\!N}$ is larger than predicted. Because the $\Lambda(1405)$ resonance is considered to be the $\bar{K}\!N$ quasi-bound state, the dominant decay branch of $\bar{K}\!N\!N$ is naively expected to be $\bar{K}\!N\!N\to\pi\Sigma N$, similar to the $\Lambda(1405)\to\pi\Sigma$ decay (100%). We would then expect $\Gamma_{\bar{K}\!N\!N}$ to be similar to the decay width of $\Lambda(1405)$. However, the obtained $\Gamma_{\bar{K}\!N\!N}$ is approximately twice as large as $\Gamma_{\Lambda(1405)}$.
@@
Two reasonable scenarios can explain the large $\Gamma_{\bar{K}\!N\!N}$. First, the $\bar{K}\!N\!N$ might decay into non-mesonic $Y\!N$ channels with strength similar to the $\pi\Sigma N$ channel. This non-mesonic decay occurs via $\bar{K}$ absorption by two nucleons in the nucleus, which most theoretical calculations have not considered. Second, the $I_{\bar{K}\!N}=1$ interaction might play an important role, similar to the $I_{\bar{K}\!N}=0$ interaction, at least in the absorption channels (i.e., $\bar{K}\!N\!N\to[\pi Y]_{I=1}N$), in contrast to the theoretical predictions [10, 11]. These two decay modes cannot occur in the decay of $\Lambda(1405)$ and newly open up in the $\bar{K}\!N\!N$ three-body system. Therefore, to clarify why $\Gamma_{\bar{K}\!N\!N}$ is much larger than $\Lambda(1405)$, studying the mesonic decay modes is critical.
@@
To investigate the mesonic decay branches of $\bar{K}\!N\!N$, we analyzed $K^{-}+^{3}{\rm He}$ reactions resulting in $\pi Y\!N\!N$ mesonic final states using data from the J-PARC E15 experiment. The measured final states in this study are $\pi^{-}\Lambda pp$, $\pi^{+}\Lambda nn$, and $\pi^{\mp}\Sigma^{\pm}pn$. As described in Ref. [2], the $\bar{K}\!N\!N$ state is produced through the $\bar{K}$ and $N$ substitution via a nucleon knockout reaction on ${{}^{3}\rm He}$. Therefore, we focus on the $\pi Y\!N\!N$ channels that originate from the reaction processes
@@
where $N^{\prime}$ is emitted in the forward direction. The backscattered intermediate $\bar{K}$ is subsequently absorbed by the two spectator nucleons as
@@
where $\bar{K}\!N\!N$ production could contribute. The relevant $K^{-}+{{}^{3}\rm He}$ reaction channels of interest can be summarized as
@@
where $\{\pi Y\!N\}$ is the particle set to be investigated and $N^{\prime}$ is the emitted nucleon in the primary reaction (Eq. (1)). Each final state is reconstructed as described on the right side of the equations, where $[\pi(\pi N)N]_{\rm det}$ represents detected particles and $N_{\rm miss}$ is an undetected particle identified using the missing mass method. The $\bar{K}\!N\!N$ signal is expected to be observed in the $[\pi Y\!N]$ invariant-mass spectra for the $I_{3}=-1/2$ state in Eq. (3), and the $I_{3}=+1/2$ state in Eqs. (4) and (5). Following the analysis in Ref. [2], we performed a two-dimensional analysis on the $[\pi Y\!N]$ invariant mass and the momentum transfer to the $[\pi Y\!N]$ system to isolate the signal of $\bar{K}\!N\!N$.
These reaction processes can couple to $Y^{*}$ resonances. In particular, two $Y^{*}$ resonances below the $\bar{K}\!N$ mass threshold, $\Lambda(1405)$ and $\Sigma(1385)$, would importantly contribute to the $\bar{K}$ absorption inside $\bar{K}\!N\!N$. Therefore, we analyzed the invariant-mass spectra of the $[\pi Y]$ subsystem to study the contributions of these $Y^{*}$ resonances.
This article is organized as follows. In Sec. II, we briefly introduce the J-PARC E15 experiment and describe the essential components of the detector system. In Sec. III, we present the analysis procedure used to select the $\pi Y\!N\!N$ channels. The obtained spectra are shown in the results in Sec. IV. In Sec. V, we perform a model fitting using the functions introduced in Ref. [2]. Specifically, we discuss the $\bar{K}\!N\!N$ production and the quasi-free $\bar{K}$ absorption processes. In the fitting procedure, we consider a background originating from the misidentification of $N^{\prime}$ and $N$ in Eqs. (1) and (2), respectively. Finally, we compare the cross-sections of $K^{-}+{{}^{3}\rm He}$ reactions leading to $\pi Y\!N+N^{\prime}$ channels with the non-mesonic $\Lambda p+n$ channel.
@@ II Experiment
We briefly overview the experimental setup of J-PARC E15, focusing on the key apparatus used to analyze the $\pi Y\!N\!N$ channels. The experimental configuration has been described in detail in the literature [12, 13, 14]. We took the current dataset in parallel to that used for the $K^{-}+\,^{3}{\rm He}\to\Lambda p+n$ analysis [2] with the same trigger condition. In the experiment, a focused $K^{-}$ beam interacted with a liquid 3He target. The particles produced in the $K^{-}+{{}^{3}\rm He}$ reaction were detected by a cylindrical detector system (CDS) surrounding the target. The CDS consists of a cylindrical drift chamber (CDC) and a cylindrical hodoscope counter made of a plastic scintillator (CDH).
The CDS operates in a magnetic field of $0.7~{}{\rm T}$ along the beam axis ($z$-axis). The tracking information provided by the CDC enables the analysis of charged particle momenta. We determined the starting point of the charged particle or the reaction point from the CDC track and $K^{-}$ beam track by searching for the closest point of the two tracks. In addition, particle identification for each track is performed using momentum and time-of-flight information obtained from the CDH.
For neutron detection, we used the CDH, which consists of plastic scintillators measuring 30 mm in thickness and 790 mm in length. The inner surface of the CDH was positioned at a radius of 544 mm from the beam axis. Using a Monte Carlo simulation based on Geant4, we estimated the typical neutron detection efficiency to be 3–9% (at neutron momenta of 1–0.2 ${\rm GeV}/c$). The neutron momentum was determined using the time-of-flight information and neutron flight length obtained from the reaction point, and the $z$-position along the CDH deduced by the time difference between two PMTs at both ends.
@@ III Analysis
We focused on the following four channels: ${\pi^{-}\Lambda p}+p^{\prime}$ (Eq. (3)), ${\pi^{+}\Lambda n}+n^{\prime}$ (Eq. (4)), and ${\pi^{\mp}\Sigma^{\pm}p}+n^{\prime}$ (Eq. (5)). Among these four channels, only the ${\pi^{-}\Lambda p}+p^{\prime}$ channel can be measured without detecting a neutron; specifically, we required the detection of $(pp\pi^{-}\pi^{-})$ in the CDS. In the other three channels, neutron detection is necessary, along with detecting the remaining three charged particles ($p\pi^{+}\pi^{-}$).
The CDS detects the ${\pi Y\!N}$ particles in each final state. The undetected particle $N^{\prime}$ is identified through the missing mass method using the measured $K^{-}$ beam and the information from the ${\pi Y\!N}$ particles. Other ${\pi Y\!N}+N^{\prime}$ channels, including multi-neutrons or $\gamma$ in the final decay products, were omitted from the present analysis because the final states cannot be identified unambiguously.
@@ III.1 Neutron analysis
We applied the following criteria to identify neutrons. First, we required no CDC track on the fired CDH segment and both sides to eliminate fake neutron signals from charged particles impacting the material close to the CDH segment. Second, we required the energy deposited by the CDH-hit to be greater than $2~{}{\rm MeV}ee$ (electron-equivalent) to remove the accidental background recorded in the CDH. Finally, $1/\beta$ was calculated for the CDH-hit using the time-of-flight information, and events with $1.51/\beta<$ 1.5 were rejected to eliminate $\gamma$-ray background originating from the $K^{-}+{{{}^{3}}{\rm He}}$ reaction, which peaks at $1/\beta\approx 1$.
Figure 1: Mass distributions of missing particle $X$ in the ($pn\pi^{+}\pi^{-}+X$) event selection: (a) distribution before background subtraction and (b) distribution after background subtraction.
Even after implementing the aforementioned background rejection procedures, our data still contain numerous neutron-like fake signals in the neutron’s missing mass spectrum. Figure 1(a) illustrates the distribution of missing mass ($m_{X}$) in the $K^{-}+{{{}^{3}}{\rm He}}\to pn\pi^{+}\pi^{-}+X$ events. The figure shows that the signal-to-noise ratio is approximately 1:1, necessitating a thorough evaluation of the background effect to achieve a detailed analysis.
The fake neutron signals are generated through multiple reactions or scattering, reach the CDH, and are identified as neutral particles. We therefore evaluated the effect of these fake signals by simulating random CDH hits using an event-mixing method. In this method, we disconnected the CDH-hit information from other charged-particle-hit information and randomly recombined the two datasets. The gray hatched curve in the figure represents the result obtained when this event-mixing method was applied. We determined a scaling factor for the background distribution by fitting the $m_{X}$ distribution below $0.8~{}{\rm GeV}/c^{2}$, where the data should only include neutron contamination events.
Figure 1(b) shows that the background subtraction was successful in the $m_{X}=m_{n}$ region and that the $[pn\pi^{+}\pi^{-}]+n_{\rm miss}$ final state clearly remains as a peak in the missing-mass spectrum. This background subtraction procedure also agrees in the higher-mass region up to 1.2 GeV/$c^{2}$, including the $m_{X}\sim m_{\Lambda}$ region. In fact, $\Lambda$ missing events, specifically $[pn\pi^{+}\pi^{-}]+\Lambda_{\rm miss}$ events, also remain clearly visible. This agreement demonstrates the validity of the fake-neutron subtraction procedure. We evaluated the fitting error of the scaling factor, as indicated by the gray box in the figure, and considered this error to be a systematic error in the subsequent analysis.